For the following exercises, determine whether there is a minimum or maximum value to each quadratic function. Find the value and the axis of symmetry.

19. f(x) = __1 2 x2 + 3x + 1

20. f(x) = −__1 3 x2 − 2x + 3
Answer #1

A quedratc equehon ot the form fo) = ax2+bx + C %3D has maximum value when a $Is negahve a is Posih ve value wh 2) bas minimaxis of Symettny s X = -0.125 20) fes = - 13x2 -ax+3 a is negative so we will have maximum at r= =b ఏం a= -13 b=-2, (=3 %3D

if there is any error in given expression .and given expression are as follows ,then answer are given below

19) if f(x)=x2+3x+1
minimum value = -1.25

axis of symmetry is x=-1.5

20) if f(x)=-x2-2x+3

maximum value =4

axis of symmetry x=-1

If there is no error ,then use answers in image